OpenStudy (anonymous):

how many license plates are possible if four letters are followed by two digits?

OpenStudy (anonymous):

Hint: Permutation problem.

OpenStudy (anonymous):

whats that?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 26 \times 26 \times 26 \times 26 \times 10 \times 10 }\) Because there are 26 letters and 10 digits.

Parth (parthkohli):

Or if they don't repeat: \(26 \times 25 \times 24 \times 23 \times 10 \times 9\)

OpenStudy (anonymous):

counting principle 26 possible letters in each of four slots \(26\times 26\times 26\times 26=26^4\) possibilities for the letters then 10 possible digits in each of 2 slots so \(10\times 10=10^2\) possibilities for the letters all together get what parthkohli said

OpenStudy (anonymous):

so the answer is 457,076?

OpenStudy (anonymous):

are you sure?